OpenStudy (anonymous):

Jerome walked 2/3 mile in the morning and 11/12 mile in the evening. How many miles did he walk in all?

OpenStudy (anonymous):

2/3 + 11/12 To add these fractions you need to make their denominators the same.

OpenStudy (anonymous):

Ok how ;c?

OpenStudy (texaschic101):

look at your denominators....3 and 12...now what do both of these numbers both go into evenly ?

OpenStudy (anonymous):

6

OpenStudy (texaschic101):

does 12 go into 6 evenly ? No. Its not what goes into them equally, it is what they go into evenly. 3 -- 3,6,9,12,15 12 -- 12,24,36 now what number do they have in common ?

OpenStudy (anonymous):

12

OpenStudy (texaschic101):

BINGO....that is your common denominator. Now we want to make both fractions have the same denominator of 12. 2 -- = --- 3 12 to find the numerator : 3 goes into 12 four times.....4 x 2 = 8...so your new fraction is 8/12. Do you understand what I just did ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

I have a problem

OpenStudy (anonymous):

Theres no 8/12 as a answer though

OpenStudy (anonymous):

http://prntscr.com/3kmrxz

OpenStudy (texaschic101):

11 --- = --- 12 12 now...12 goes into 12 one time. 1 x 11 = 11. This fraction becomes 11/12 so now we have : 2/3 = 8/12 11/12 = 11/12 now we can add them because they have the same denominator. 8/12 + 11/12 = 19/12 or 1 7/12

OpenStudy (anonymous):

OHH

OpenStudy (texaschic101):

do you know how I got 1 7/12 from 19/12 ?

OpenStudy (anonymous):

Yes

OpenStudy (texaschic101):

good....1 7/12 is your answer

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

I have two more

OpenStudy (texaschic101):

lets see them

OpenStudy (anonymous):

The base of a triangle is 5/8 feet. The height is 1/3 feet. How much greater is the base of the triangle than the height?

OpenStudy (texaschic101):

so the problem is 5/8 - 1/3 common denominator of 8 and 3....do you know it ?

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

Wait no

OpenStudy (texaschic101):

3 -- 3,6,9,12,15,18,21,24,27 8 -- 8,16,24,32,40 what is your common number ?

OpenStudy (anonymous):

24

OpenStudy (texaschic101):

yep...now we have to convert the fraction to where both denominators are 24. I am sorry but I am going to have to go, so I will do this one for you 5/8 = 15/24 1/3 = 8/24 15/24 - 8/24 = 7/24 I am sorry I cant stay and help more...I will be back on later

OpenStudy (anonymous):

Ok